OpenStudy (anonymous):

Layna can paint a 350-square-foot room in 4 hours. It takes Bebe 7 hours to paint the same room. Which equation can be used to find t, the time it takes Layna and Bebe to paint the room together?

OpenStudy (anonymous):

Add them. Divide by 2. that should be the awnser so... 5.5 hours?

OpenStudy (anonymous):

look if you want to find the time of working 2 person together use this : if time for person 1 = t1 time for person 2 = t2 We have : 1/t1+1/t2=x and then have 1/x Got it ?:)

OpenStudy (anonymous):

thank you ali

OpenStudy (anonymous):

So...that would still be 5.5 hours? this would help me...

OpenStudy (anonymous):

For example : if : person 1 = 2min person 2 = 4 min person 2 and one together : 1/2+1/4=2+1/4 = 3/4 then x = 4/3 and have : 4/3=x/60 x=80 s Got it ?:)

OpenStudy (anonymous):

the options are 1. 1/4t + 1/7t = 1 2. 1/4t * 1/7t = 1 3. 4t + 7t = 1 4. 4t - 7t = 1

OpenStudy (anonymous):

*derp face* Nope but i kind of... So your saying if i took 3 hours to rake a yard alone...and another person takes 5 hours to rake the same yard...then...take...3 and break it in half to get 1.5 and 5 in half for 2.5 and then add them together to get 4 hours? but that would be the same thing i did before correct...so no...i guess i dont get it hu...

OpenStudy (anonymous):

Look : In there we have : t1=1/4 t2=1/7 and we have : 1/4t+1/7t

OpenStudy (anonymous):

thank you again E.ali

OpenStudy (anonymous):

Nope still dont get it but its okay...thanks for trying...Ill probably learn it later...and be like Wtf is this crap? then ill remember that someone on the interwebs said it and ill be like OH CRAP THATS IT!